# Parrondo’s paradox in susceptible-infectious-susceptible dynamics over periodic temporal networks

### v1

@@ 3.2 Relationships between the amount of interaction and epidemic spreading
We propose that an advantage of the Parrondo’s paradox, or using periodic switching networks showing it, is to have comparably less epidemic spreading for the given total amount of interaction between individuals. To examine this point, we define the total amount of interaction between individuals in the case of static networks by the sum of all the entries of the adjacency matrix. In the case of undirected networks, this definition double counts the off-diagonal entries but provides a correctly normalized total amount of interaction. Imagine a static SBM with two blocks in which each block has $N/2$ individuals and the edge exists between two individuals in the first subpopulation with probability $a_{11}/N$, for example. Then, the expected number of edges, which is equivalent to the amount of interaction in the case of static networks, within the first subpopulation is $\frac{a_{11}}{N}\times\left[\frac{1}{2}\times\frac{N}{2}\left(\frac{N}{2}-1\right)\right]\approx\frac{Na_{11}}{8}$. The expected number of edges between the two subpopulations is $\frac{a_{12}}{N}\times\left(\frac{N}{2}\right)^{2}=\frac{Na_{12}}{4}$, justifying the double counting. For periodic switching networks, we further take the time average over one cycle of the total amount of interaction. The total amount of interaction for periodic switching networks is equal to $\sum_{\ell^{\prime}=1}^{\ell}\tau_{\ell^{\prime}}\sum_{i=1}^{N}\sum_{j=1}^{N}a_{ij}^{(\ell^{\prime})}/T$.
@@
Figure 3: Leading eigenvalues as a function of the total amount of interaction for a two-node periodic switching network and the constituent static networks. The value of $\mu$ $(=0.5)$, $\textbf{A}^{(1)}$, $\textbf{A}^{(2)}$, and the switching schedule (i.e., $\ell=2$, $T=7$, and $\tau_{1}=5$, and hence $\tau_{2}=T-\tau_{1}=2$) are the same as those used in Fig. 2. The shaded region represents the region of “$\beta\times(\text{amount of interaction})$” in which the Parrondo’s paradox occurs.
3.3 Generality of the Parrondo’s paradox
@@
Next we examine the generality of the Parrondo’s paradox. First, we analyze some other two-node periodic switching networks. We use the same switching schedule (i.e., $\ell=2$ alternating static networks, period $T=7$, the duration of the first static network $\tau_{1}=5$, and hence $\tau_{2}=T-\tau_{1}=2$) in the remainder of this paper including here. We consider three one-parameter families of two-node periodic switching networks and calculated the epidemic threshold for each periodic switching network. We show in Fig. 4(a)–(c) the epidemic threshold for the periodic switching network (i.e., $\beta_{\text{F}}^{*}$), first static network (i.e., $\beta_{1}^{*}$), and second static network (i.e., $\beta_{2}^{*}$). We set $\textbf{A}^{(1)}=\begin{pmatrix}\gamma&5\\ 5&0\end{pmatrix}$ and $\textbf{A}^{(2)}=\begin{pmatrix}0&5\\ 5&15\end{pmatrix}$ in Fig. 4(a), $\textbf{A}^{(1)}=\begin{pmatrix}17&5\\ 5&\gamma\end{pmatrix}$ and $\textbf{A}^{(2)}=\begin{pmatrix}\gamma&5\\ 5&15\end{pmatrix}$ in Fig. 4(b), and $\textbf{A}^{(1)}\begin{pmatrix}17&\gamma\\ \gamma&0\end{pmatrix}$ and $\textbf{A}^{(2)}=\begin{pmatrix}0&\gamma\\ \gamma&15\end{pmatrix}$ in Fig. 4(c). The shaded regions in the figure are the ranges of $\gamma$ and $\beta$ for which the Parrondo’s paradox occurs, i.e., where $\lambda_{\text{F}}<0$, $\lambda_{\max}(\textbf{M}^{(1)})>0$, and $\lambda_{\max}(\textbf{M}^{(2)})>0$ are simultaneously satisfied. We find that the paradox occurs for some ranges of $\gamma$.
@@
To verify that the Parrondo’s paradox can also be present in large networks, we similarly calculated the leading eigenvalues for networks with $N=2000$ nodes generated by an SBM that creates large networks with community structure as discussed in Appendix B. We define the SBMs to have two communities, each with $N/2=1000$ nodes, and obtain their edge probabilities using the entries in the adjacency matrices of the three families of $2\times 2$ periodic switching networks used in Fig. 4(a)–(c). Summing the entries across SBM blocks recovers the original $2\times 2$ matrices, in expectation. We show in Fig. 4(d)–(f) the epidemic thresholds and the region of pairs of $\gamma$ and $\beta$ in which the Parrondo’s paradox is present for the networks with $N=2000$ nodes generated by the SBM. The results for the SBM are close to those for the corresponding two-node networks shown in Fig. 4(a)–(c).
@@
Figure 4: Parrondo’s paradox over other periodically switching networks with two nodes. We show the epidemic threshold for the periodic switching network (i.e., $\beta_{\text{F}}^{*}$; black lines), $\textbf{A}^{(1)}$ (i.e., $\beta_{1}^{*}$; blue lines), and $\textbf{A}^{(2)}$ (i.e., $\beta_{2}^{*}$; magenta lines). The shaded region shows the values of $\gamma$ and $\beta$ for which the paradox occurs. (a)–(c): Two-node networks. (d)–(f): Networks with $N=2000$ nodes and two communities generated by the SBM. (a) and (d): $\textbf{A}^{(1)}=\begin{pmatrix}\gamma&5\\ 5&0\end{pmatrix}$, $\textbf{A}^{(2)}=\begin{pmatrix}0&5\\ 5&15\end{pmatrix}$. (b) and (e): $\textbf{A}^{(1)}=\begin{pmatrix}17&5\\ 5&\gamma\end{pmatrix}$, $\textbf{A}^{(2)}=\begin{pmatrix}\gamma&5\\ 5&15\end{pmatrix}$. (c) and (f): $\textbf{A}^{(1)}=\begin{pmatrix}17&\gamma\\ \gamma&0\end{pmatrix}$, $\textbf{A}^{(2)}=\begin{pmatrix}0&\gamma\\ \gamma&15\end{pmatrix}$.
@@
We defined the Parrondo’s paradox as a situation in which $\lambda_{\text{F}}<0$, $\lambda_{\max}(\textbf{M}^{(1)})>0$, and $\lambda_{\max}(\textbf{M}^{(2)})>0$ simultaneously hold true. Because these leading eigenvalues monotonically increase as $\beta$ increases, a periodic switching network shows the Parrondo’s paradox for some values of $\beta$ if and only if $\beta_{\text{F}}^{*}>\max\{\beta_{1}^{*},\beta_{2}^{*}\}$, as shown by the shaded regions in Figs. 2 and 4. We generated $10^{4}$ independent pairs of $\textbf{A}^{(1)}$ and $\textbf{A}^{(2)}$ to find that the paradox is present in 28.7% of the periodic switching networks. Therefore, we conclude that the Parrondo’s paradox is not a rare phenomenon, at least for $2\times 2$ adjacency matrices. We did not find any periodic switching network that showed the opposite paradoxical behavior, i.e., $\beta_{\text{F}}^{*}<\min\{\beta_{1}^{*},\beta_{2}^{*}\}$.
@@
As our final set of experiments to investigate the generality of the Parrondo’s paradox, we study how it is affected by the number of node. We first examine the possibility of the paradox with $3\times 3$ periodic switching networks. Let us consider arbitrarily selected one-parameter families of three-node periodic switching networks. The epidemic thresholds for the periodic switching network (i.e., $\beta_{\text{F}}^{*}$) and those for the static networks (i.e., $\beta_{1}^{*}$, $\beta_{2}^{*}$) are shown in Fig. 5(a) for $\textbf{A}^{(1)}=\begin{pmatrix}\gamma&5&5\\ 5&0&5\\ 5&5&15\end{pmatrix}$ and $\textbf{A}^{(2)}=\begin{pmatrix}14&5&5\\ 5&0&5\\ 5&5&\gamma\end{pmatrix}$ and Fig. 5(b) for $\textbf{A}^{(1)}=\begin{pmatrix}24&\gamma&5\\ 5&0&\gamma\\ \gamma&5&15\end{pmatrix}$ and $\textbf{A}^{(2)}=\begin{pmatrix}14&\gamma&5\\ 5&0&\gamma\\ \gamma&5&24\end{pmatrix}$. The shaded regions represent the region of $\gamma$ and $\beta$ values in which the Parrondo’s paradox occurs. We observe that the paradox occurs for a range of $\gamma$ in both families of periodic switching networks. For the periodic switching network shown in Fig. 5(a), the ranges of $\gamma$ for which the paradox occurs are disjoint; the paradox occurs for small and large $\gamma$ values, but not for intermediate $\gamma$ values.
@@
Figure 5: Epidemic thresholds for periodic switching networks and static networks with three nodes. (a): $\textbf{A}^{(1)}=\begin{pmatrix}\gamma&5&5\\ 5&0&5\\ 5&5&15\end{pmatrix}$ and $\textbf{A}^{(2)}=\begin{pmatrix}14&5&5\\ 5&0&5\\ 5&5&\gamma\end{pmatrix}$. (b): $\textbf{A}^{(1)}=\begin{pmatrix}24&\gamma&5\\ 5&0&\gamma\\ \gamma&5&15\end{pmatrix}$ and $\textbf{A}^{(2)}=\begin{pmatrix}14&\gamma&5\\ 5&0&\gamma\\ \gamma&5&24\end{pmatrix}$. See the caption of Fig. 4 for the legends. The shaded regions represent the values of $\gamma$ and $\beta$ for which the Parrondo’s paradox occurs. In (b), the blue line is not visible because it is almost completely hidden behind the magenta line.
@@
To further assess the generality of the Parrondo’s paradox for periodic switching networks with three nodes, we retain the same switching schedule for simplicity and generate symmetric adjacency matrices parameterized as $\textbf{A}^{(1)}=\begin{pmatrix}\gamma_{1}&\gamma_{2}&\gamma_{3}\\ \gamma_{2}&\gamma_{4}&\gamma_{5}\\ \gamma_{3}&\gamma_{5}&\gamma_{6}\end{pmatrix}$ and $\textbf{A}^{(2)}=\begin{pmatrix}\gamma_{7}&\gamma_{8}&\gamma_{9}\\ \gamma_{8}&\gamma_{10}&\gamma_{11}\\ \gamma_{9}&\gamma_{11}&\gamma_{12}\end{pmatrix}$. We set $\gamma_{1}=1$ without loss of generality. Then, we independently draw $\gamma_{2}$, $\ldots$, $\gamma_{12}$ from the uniform density on $[0,5]$. We find that, among $10^{4}$ pairs of the generated $\textbf{A}^{(1)}$ and $\textbf{A}^{(2)}$, the paradox is present, with $\beta_{\text{F}}^{*}>\max\{\beta_{1}^{*},\beta_{2}^{*}\}$, in 12.0% of the periodic switching networks, and that there is no case with $\beta_{\text{F}}^{*}<\min\{\beta_{1}^{*},\beta_{2}^{*}\}$.
@@
3.4 Relationships between the Parrondo’s paradox and anti-phase oscillations
To better understand mechanisms of the Parrondo’s paradox, we carried out additional experiments motivated by the following observation. Figure 2 suggests that the Parrondo’s paradox may be characterized by an anti-phase oscillation between $x_{1}(t)$ and $x_{2}(t)$. However, this association is not straightforward already with Fig. 2 because the figure shows anti-phase oscillations for a $\beta$ value at which the paradox occurs (see Fig. 2(d)) as well as for a $\beta$ value at which it does not (see Fig. 2(c)). Nevertheless, the anti-phase oscillation is absent for $\beta$ values that are far from where the paradox is present (see Fig. 2(b) and 2(e)).
We measure the extent of anti-phase oscillation by the fraction of time, $q$, during one period in which $\text{d}x_{1}(t)/\text{d}t$ and $\text{d}x_{2}(t)/\text{d}t$ have the opposite sign. We measure $q$ in the fifth cycle, i.e., using $t\in[28,35]$, and use the initial condition $(x_{1}(0),x_{2}(0))=(0.1,0.1)$. Consider the $10^{4}$ periodic switching networks with two communities used in section 3.3. We recall that $28.7\%$ of the periodic switching networks (i.e., 2,870 networks) show the Parrondo’s paradox. For each of these 2,870 networks, we select the middle value of the range of $\beta$ in which the paradox occurs and compute $q$ at the selected $\beta$ value. At this value of $\beta$, perfect anti-phase synchronization, i.e., $q=1$, occurs for 2867 ($99.9\%$) of the 2870 networks. For comparison, we have also computed $q$ for the rest of the networks (i.e., 7,130 out of the $10^{4}$ networks). We find that the maximum, minimum, mean and standard deviation of $q$ for these 7,130 networks is $0.974$, 00, $0.366$, and $0.305$, respectively. Therefore, although some networks show strong anti-phase oscillation despite the lack of the Parrondo’s paradox behavior, most networks not showing the paradox have substantially smaller $q$ values than the networks showing the paradox.
We show examples of anti-phase dynamics in Appendix C. It should be noted that $q=1$ holds true for both inside and outside the range of $\beta$ in which the paradox is present for some networks. In our examples, $q$ is much larger for the periodic switching networks showing the paradox than those not showing the paradox. We also observe some anti-phase behavior in the four periodic switching networks that do not show the Parrondo’s paradox for any value of $\beta$, but to a limited extent (i.e., $q\leq 0.7$). These results are consistent with the population results with $10^{4}$ networks described above. In sum, we conclude that anti-phase oscillation occurs more frequently in the presence or approximate presence of the Parrondo’s paradox than otherwise.

### v2

@@ 3.2 Relationships between the amount of interaction and epidemic spreading
We propose that an advantage of the Parrondo paradox, or using periodic switching networks showing it, is to have comparably less epidemic spreading for the given total amount of interaction between individuals. To examine this point, we define the total amount of interaction between individuals in the case of static networks by the sum of all the entries of the adjacency matrix. In the case of undirected networks, this definition double counts the off-diagonal entries but provides a correctly normalized total amount of interaction. Imagine a static SBM with two blocks in which each block has $N/2$ individuals and the edge exists between two individuals in the first subpopulation with probability $a_{11}/N$, for example. Then, the expected number of edges, which is equivalent to the amount of interaction in the case of static networks, within the first subpopulation is $\frac{a_{11}}{N}\times\left[\frac{1}{2}\times\frac{N}{2}\left(\frac{N}{2}-1\right)\right]\approx\frac{Na_{11}}{8}$. The expected number of edges between the two subpopulations is $\frac{a_{12}}{N}\times\left(\frac{N}{2}\right)^{2}=\frac{Na_{12}}{4}$, justifying the double counting. For periodic switching networks, we further take the time average over one cycle of the total amount of interaction. The total amount of interaction for periodic switching networks is equal to $\sum_{\ell^{\prime}=1}^{\ell}\tau_{\ell^{\prime}}\sum_{i=1}^{N}\sum_{j=1}^{N}a_{ij}^{(\ell^{\prime})}/T$.
@@
Figure 4: Leading eigenvalues as a function of the total amount of interaction for a two-node periodic switching network and the constituent static networks. The value of $\mu$ $(=0.5)$, $\textbf{A}^{(1)}$, $\textbf{A}^{(2)}$, and the switching schedule (i.e., $\ell=2$, $T=7$, and $\tau_{1}=5$, and hence $\tau_{2}=T-\tau_{1}=2$) are the same as those used in Fig. 3. The shaded region represents the region of “$\beta\times(\text{amount of interaction})$” in which the Parrondo paradox occurs.
3.3 Generality of the Parrondo paradox
@@
Next we examine the generality of the Parrondo paradox with respect to the connection strength between and within subpopulations, the number of subpopulations up to $N=5$, and larger networks of individuals (as opposed to networks of subpopulations) with distinct two subpopulations. First, we analyze some other two-node periodic switching networks. We use the same switching schedule (i.e., $\ell=2$ alternating static networks, period $T=7$, the duration of the first static network $\tau_{1}=5$, and hence $\tau_{2}=T-\tau_{1}=2$) in the remainder of this paper including here. We consider three one-parameter families of two-node periodic switching networks and calculated the epidemic threshold for each periodic switching network. We show in Fig. 5(a)–(c) the epidemic threshold for the periodic switching network (i.e., $\beta_{\text{F}}^{*}$), first static network (i.e., $\beta_{1}^{*}$), and second static network (i.e., $\beta_{2}^{*}$). We set $\textbf{A}^{(1)}=\begin{pmatrix}\gamma&5\\ 5&0\end{pmatrix}$ and $\textbf{A}^{(2)}=\begin{pmatrix}0&5\\ 5&15\end{pmatrix}$ in Fig. 5(a), $\textbf{A}^{(1)}=\begin{pmatrix}17&5\\ 5&\gamma\end{pmatrix}$ and $\textbf{A}^{(2)}=\begin{pmatrix}\gamma&5\\ 5&15\end{pmatrix}$ in Fig. 5(b), and $\textbf{A}^{(1)}\begin{pmatrix}17&\gamma\\ \gamma&0\end{pmatrix}$ and $\textbf{A}^{(2)}=\begin{pmatrix}0&\gamma\\ \gamma&15\end{pmatrix}$ in Fig. 5(c). The shaded regions in the figure are the ranges of $\gamma$ and $\beta$ for which the Parrondo paradox occurs, i.e., where $\lambda_{\text{F}}<0$, $\lambda_{\max}(\textbf{M}^{(1)})>0$, and $\lambda_{\max}(\textbf{M}^{(2)})>0$ are simultaneously satisfied. We find that the paradox occurs for some ranges of $\gamma$.
@@
To verify that the Parrondo paradox can also be present in large networks, we similarly calculated the leading eigenvalues for networks with $N=2000$ nodes generated by an SBM that creates large networks with community structure as discussed in Appendix C. We define the SBMs to have two communities, each with $N/2=1000$ nodes, and obtain their edge probabilities using the entries in the adjacency matrices of the three families of $2\times 2$ periodic switching networks used in Fig. 5(a)–(c). Summing the entries across SBM blocks recovers the original $2\times 2$ matrices, in expectation. We show in Fig. 5(d)–(f) the epidemic thresholds and the region of pairs of $\gamma$ and $\beta$ in which the Parrondo paradox is present for the networks with $N=2000$ nodes generated by the SBM. The results for the SBM are close to those for the corresponding two-node networks shown in Fig. 5(a)–(c).
@@
Figure 5: Parrondo’s paradox over other periodically switching networks with two nodes. We show the epidemic threshold for the periodic switching network (i.e., $\beta_{\text{F}}^{*}$; black lines), $\textbf{A}^{(1)}$ (i.e., $\beta_{1}^{*}$; blue lines), and $\textbf{A}^{(2)}$ (i.e., $\beta_{2}^{*}$; magenta lines). The shaded region shows the values of $\gamma$ and $\beta$ for which the paradox occurs. (a)–(c): Two-node networks. (d)–(f): Networks with $N=2000$ nodes and two communities generated by the SBM. Panels (a) and (d) use the same adjacency matrices $\textbf{A}^{(1)}$ and $\textbf{A}^{(2)}$, indicated below (d). Same for (b) and (e), and for (c) and (f).
@@
We defined the Parrondo paradox as a situation in which $\lambda_{\text{F}}<0$, $\lambda_{\max}(\textbf{M}^{(1)})>0$, and $\lambda_{\max}(\textbf{M}^{(2)})>0$ simultaneously hold true. Because these leading eigenvalues monotonically increase as $\beta$ increases, a periodic switching network shows the Parrondo paradox for some values of $\beta$ if and only if $\beta_{\text{F}}^{*}>\max\{\beta_{1}^{*},\beta_{2}^{*}\}$, as shown by the shaded regions in Figs. 3 and 5. We generated $10^{4}$ independent pairs of $\textbf{A}^{(1)}$ and $\textbf{A}^{(2)}$ to find that the paradox is present in 28.7% of the periodic switching networks. Therefore, we conclude that the Parrondo paradox is not a rare phenomenon, at least for $2\times 2$ adjacency matrices. We did not find any periodic switching network that showed the opposite paradoxical behavior, i.e., $\beta_{\text{F}}^{*}<\min\{\beta_{1}^{*},\beta_{2}^{*}\}$.
@@
As our final set of experiments to investigate the generality of the Parrondo paradox, we study how it is affected by the number of node. We first examine the possibility of the paradox with $3\times 3$ periodic switching networks. Let us consider arbitrarily selected one-parameter families of three-node periodic switching networks. The epidemic thresholds for the periodic switching network (i.e., $\beta_{\text{F}}^{*}$) and those for the static networks (i.e., $\beta_{1}^{*}$, $\beta_{2}^{*}$) are shown in Fig. 6(a) for $\textbf{A}^{(1)}=\begin{pmatrix}\gamma&5&5\\ 5&0&5\\ 5&5&15\end{pmatrix}$ and $\textbf{A}^{(2)}=\begin{pmatrix}14&5&5\\ 5&0&5\\ 5&5&\gamma\end{pmatrix}$ and Fig. 6(b) for $\textbf{A}^{(1)}=\begin{pmatrix}24&\gamma&5\\ 5&0&\gamma\\ \gamma&5&15\end{pmatrix}$ and $\textbf{A}^{(2)}=\begin{pmatrix}14&\gamma&5\\ 5&0&\gamma\\ \gamma&5&24\end{pmatrix}$. The shaded regions represent the region of $\gamma$ and $\beta$ values in which the Parrondo paradox occurs. We observe that the paradox occurs for a range of $\gamma$ in both families of periodic switching networks. For the periodic switching network shown in Fig. 6(a), the ranges of $\gamma$ for which the paradox occurs are disjoint; the paradox occurs for small and large $\gamma$ values, but not for intermediate $\gamma$ values.
@@
Figure 6: Epidemic thresholds for periodic switching networks and static networks with three nodes. See the caption of Fig. 5 for the legends. The shaded regions represent the values of $\gamma$ and $\beta$ for which the Parrondo paradox occurs. In (b), the blue line is not visible because it is almost completely hidden behind the magenta line.
@@
To further assess the generality of the Parrondo paradox for periodic switching networks with three nodes, we retain the same switching schedule for simplicity and generate symmetric adjacency matrices parameterized as $\textbf{A}^{(1)}=\begin{pmatrix}\gamma_{1}&\gamma_{2}&\gamma_{3}\\ \gamma_{2}&\gamma_{4}&\gamma_{5}\\ \gamma_{3}&\gamma_{5}&\gamma_{6}\end{pmatrix}$ and $\textbf{A}^{(2)}=\begin{pmatrix}\gamma_{7}&\gamma_{8}&\gamma_{9}\\ \gamma_{8}&\gamma_{10}&\gamma_{11}\\ \gamma_{9}&\gamma_{11}&\gamma_{12}\end{pmatrix}$. We set $\gamma_{1}=1$ without loss of generality. Then, we independently draw $\gamma_{2}$, $\ldots$, $\gamma_{12}$ from the uniform density on $[0,5]$. We find that, among $10^{4}$ pairs of the generated $\textbf{A}^{(1)}$ and $\textbf{A}^{(2)}$, the paradox is present, with $\beta_{\text{F}}^{*}>\max\{\beta_{1}^{*},\beta_{2}^{*}\}$, in 12.0% of the periodic switching networks, and that there is no case with $\beta_{\text{F}}^{*}<\min\{\beta_{1}^{*},\beta_{2}^{*}\}$.
@@
3.4 Relationships between the Parrondo paradox and anti-phase oscillations
To better understand mechanisms of the Parrondo paradox, we carried out additional experiments motivated by the following observation. Figure 3 suggests that the Parrondo paradox may be characterized by an anti-phase oscillation between $x_{1}(t)$ and $x_{2}(t)$. However, this association is not straightforward already with Fig. 3 because the figure shows anti-phase oscillations for a $\beta$ value at which the paradox occurs (see Fig. 3(d)) as well as for a $\beta$ value at which it does not (see Fig. 3(c)). Nevertheless, the anti-phase oscillation is absent for $\beta$ values that are far from where the paradox is present (see Fig. 3(b) and 3(e)).
We measure the extent of anti-phase oscillation by the fraction of time, $q$, during one period in which $\text{d}x_{1}(t)/\text{d}t$ and $\text{d}x_{2}(t)/\text{d}t$ have the opposite sign. We measure $q$ in the fifth cycle, i.e., using $t\in[28,35]$, and use the initial condition $(x_{1}(0),x_{2}(0))=(0.1,0.1)$. Consider the $10^{4}$ periodic switching networks with two communities used in section 3.3. We recall that $28.7\%$ of the periodic switching networks (i.e., 2,870 networks) show the Parrondo paradox. For each of these 2,870 networks, we select the middle value of the range of $\beta$ in which the paradox occurs and compute $q$ at the selected $\beta$ value. At this value of $\beta$, perfect anti-phase synchronization, i.e., $q=1$, occurs for 2867 ($99.9\%$) of the 2870 networks. For comparison, we have also computed $q$ for the rest of the networks (i.e., 7,130 out of the $10^{4}$ networks). We find that the maximum, minimum, mean and standard deviation of $q$ for these 7,130 networks is $0.974$, 00, $0.366$, and $0.305$, respectively. Therefore, although some networks show strong anti-phase oscillation despite the lack of the Parrondo paradox behavior, most networks not showing the paradox have substantially smaller $q$ values than the networks showing the paradox.
We show examples of anti-phase dynamics in Appendix D. It should be noted that $q=1$ holds true for both inside and outside the range of $\beta$ in which the paradox is present for some networks. In our examples, $q$ is much larger for the periodic switching networks showing the paradox than those not showing the paradox. We also observe some anti-phase behavior in the four periodic switching networks that do not show the Parrondo paradox for any value of $\beta$, but to a limited extent (i.e., $q\leq 0.7$). These results are consistent with the population results with $10^{4}$ networks described above. In sum, we conclude that anti-phase oscillation occurs more frequently in the presence or approximate presence of the Parrondo paradox than otherwise.